Suppose that 20, 21, 22, ... is sequence defined as follows. do = 5,21 = 16,0 integers n > 2. Prove that an = 3.2 +2.5 for
Answer #1

90=5, a = 16 nr 2 an= 7ant-loan-2 n72 To solve this recurrence Relation put anaxh xh = 72h -1 10xn-2 - divide all by a 2²=74and a, = 16 16= A.2 t 5-13 16=2A + SB 10 = 2A + 2B 6 = 3B ---A= S-B=5-2 =3. .. an= 3.2h $2.5h an= 3.2h+2. ř is a solletion fe